Akoshia has more money than komley . If Akoshia gave komley ¢ 22, Akoshia would then have twice as much as komley. How much does each one actually have?


1
Expert's answer
2021-03-09T08:08:34-0500

Let Akoshia has ¢"x." Komley has ¢"y," "x>y."


"x-20=y+20"

"x+22=2(y-22)"

Then


"x=y+40"

"y+40+22=2y-44"

"y=106"

"x=146"


Akoshia has ¢146, Komley has ¢106.